Find the sum of the infinite series
315+189+113.4+68.04+.....

Answer by fractalier(6550)

We need the ratio r...here
r = 189/315 = .6
Then an infinite sum is found by
S = a1 / (1-r) = 315 / (1 - .6) = 315 / .4 = 787.5